MATH 565: Foundations of Mathematics II Syllabus for Spring 2015

Stephen G. Simpson Pennsylvania State University www.math.psu.edu/simpson/courses/math565

January 12, 2015

The current plan is to go through my book Subsystems of Second Order Arithmetic, starting at the beginning. Of course it will not be possible to cover the entire book. I will lend a copy of the book to each student who doesn’t already own one.

As needed, depending on the students’ background knowledge, I plan to supplement the material from my book with material from my lecture notes for MATH 457 and 557 (Mathematical Logic) and MATH 558 (Foundations of Mathematics I). These lecture notes are available at

www.math.psu.edu/simpson/notes/.

Could each student please make an individual appointment with me? The purpose of the appointment is to discuss the student’s background and interests. At the same time, the student may pick up a copy of my book if needed.

My email is simpson@math.psu.edu and my office address is 305 McAllister Building.